OpenStudy (anonymous):

Maureen cut a 20-cm ribbon into exactly three pieces. The first piece is 3 cm shorter than the second piece. The third piece is 4 cm shorter than the second piece. Find the length of the shortest piece.

OpenStudy (anonymous):

set : x= length of 2nd piece in cm x- 3= length of 1st piece in cm x-4= length of 3rd piece in cm then since you know they all came from the 20 cm ribbon you add them together and set them equal to 20 x+ x-3 +x-4 = 20 3x - 7= 20 3x= 27 x= 9 the shortest piece is the third piece which is x- 4 x-4 9-4 = 5 The length of the smallest piece is 5 cm. the other lengths are 9 cm 6 cm